Grade 12General Physics

Two mass M and m are moving in a circular path having radius R and r. If their time period is same then the ratio of angular velocity will be

RANJEET Choudhary
8 Years agoGrade 12

1 Answer

Arun
8 Years ago
Dear Ranjeet
 
The two particles m and M are moving in a circular motion with different radii  r and R..
Time period = T = 2 π / ω
=> T1/ T2 =  ω2 / ω1
=>ω2 / ω1 = 1  as  T1 = T2
 
Regards
Arun (askIITians forum expert)